# Generative-Contrastive Heterogeneous Graph Neural Network

### v1

@@ Abstract.
Heterogeneous Graphs (HGs) can effectively model complex relationships in the real world by multi-type nodes and edges. In recent years, inspired by self-supervised learning, contrastive Heterogeneous Graphs Neural Networks (HGNNs) have shown great potential by utilizing data augmentation and discriminators for downstream tasks. However, data augmentation is still limited due to the discrete and abstract nature of graphs. To tackle the above limitations, we propose a novel Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). Specifically, we first propose a heterogeneous graph generative learning enhanced contrastive paradigm. This paradigm includes: 1) A contrastive view augmentation strategy by using masked autoencoder. 2) Position-aware and semantics-aware positive sample sampling strategy for generate hard negative samples. 3) A hierarchical contrastive learning strategy for capturing local and global information. Furthermore, the hierarchical contrastive learning and sampling strategies aim to constitute an enhanced discriminator under the generative-contrastive perspective. Finally, we compare our model with seventeen baselines on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on node classification and link prediction tasks. To reproduce our work, we have open-sourced our code at https://github.com/xxx.
@@ 1. INTRODUCTION
Figure 1. Heterogeneous graph self-supervised learning with contrastive and generative-contrastive paradigms.
@@
Self-supervised learning has widely applied in the fields of computer vision (CV) (He et al., 2020b; Chen et al., 2020; Grill et al., 2020) and natural language processing (NLP) (Chowdhery et al., 2023; Devlin et al., 2018), leveraging unlabeled data to learn informative representations. Inspired by this trend, researchers begin to explore the combination of HGNNs and self-supervised learning. In particular, contrastive HGNNs (Wang et al., 2021; Zhu et al., 2022; Wang et al., 2023; Chen et al., 2023) have shown great potential because they are more suitable for downstream tasks by discarding the decoder and setting up a discriminator. Data augmentation (Wu et al., 2021; Yang et al., 2023) is crucial for improving the performance of contrastive learning on heterogeneous graphs. However, due to the discrete and abstract nature of graph data, it is difficult to provide reasonable data augmentation to leverage the advantages of contrastive learning (He et al., 2020b; Liu et al., 2021).
@@
Generative learning (Chowdhery et al., 2023; Tian et al., 2023; Hou et al., 2022; Devlin et al., 2018; Liu et al., 2021) can effectively solve the problem of data augmentation because it can reconstruct the data distribution without considering downstream tasks, allowing the each node embedding to be enhanced. This also provides a foundation for us to propose a generative learning enhanced contrastive paradigm for heterogeneous graphs. As shown in Figure 1, first, HGNNs utilize a meta-path to obtain a graph containing nodes of the same type. Contrastive learning (Wang et al., 2023; Zhu et al., 2022) generates supervised signals by aligning two similar views constructed through random edge dropping; while generative-contrastive learning (Goodfellow et al., 2020; Yang et al., 2023; Wu et al., 2023) generates fake samples by training an autoencoder and uses a discriminator to distinguish them. Despite the obvious advantages of generative-contrastive paradigm, there are still some challenges for heterogeneous graphs:
@@
CH1: Contrastive view challenge. The current augmentation strategies (Wang et al., 2023, 2021; Zhu et al., 2022; Lin et al., 2022; Wu et al., 2021), such as structural and feature augmentation, still have limitations. Structural augmentation (Wang et al., 2023; Zhu et al., 2022; Wu et al., 2021), which involves randomly dropout nodes and edges, could disrupt the graph integrity since nodes and edges are the fundamental properties of a graph. On the other hand, feature augmentation (Wang et al., 2021; Lin et al., 2022), which perturbs node attributes, ignores the unique features of each node’s input. In such methods, some nodes have more edges and embeddings than others, resulting in unbalanced supervised signals. Exploring how to better utilize meta-paths to design enhanced contrastive views remains a challenge that needs to be addressed.
@@
CH2: Sampler challenge. In heterogeneous graphs, contrastive learning (Wang et al., 2023, 2021; Zhu et al., 2022; Chen et al., 2023; Wu et al., 2021) heavy relies on high-quality negative samples. Moreover, the discriminator has stricter requirements for sampling under the generative-contrastive paradigm (Wu et al., 2023). Existing sampling strategies (Wang et al., 2023, 2021; Zhu et al., 2022) are usually based on randomness or meta-paths, which often increase false negative samples. Some recent works (Kalantidis et al., 2020; Liu et al., 2023) aim to select or synthesize hard negatives to mitigate the impact of false negatives. However, they are not suitable for heterogeneous graphs and offer limited benefits. Therefore, designing an enhanced positive sampling strategy for heterogeneous graphs is urgently needed.
@@
CH3: Contrastive mechanism challenge. Existing methods (Wang et al., 2023; Zhu et al., 2022; Tian et al., 2023) rely solely on meta-paths to capture rich semantic information in heterogeneous graphs and often overlook the influence of one-hop neighbors on target nodes. Network schema leverages direct connections between nodes to capture the local structure of target nodes. In contrast, meta-paths extract higher-order structures. Designing an effective self-supervised mechanism that considers both local and global information can better help the model work.
@@
To tackle the three challenges mentioned above, we propose a heterogeneous graph self-supervised model called Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). For CH1, GC-HGNN employs a generative masked autoencoder (MAE) to enhance contrastive views. The MAE can reconstruct node embeddings without altering the original graph structure and features. The masking strategy averages the edges of each node to ensure that each node receives the same supervised signals during the reconstruction process. For CH2, the model proposes a positive sample sampling strategy that combines location-awareness and semantic-awareness. This approach adaptively samples nodes based on the rich semantics unique to heterogeneous graphs and the global positions of nodes in the heterogeneous graph. Finally, for CH3, GC-HGNN utilizes hierarchical contrastive learning to capture one-hop and higher-order neighbors. Intra-contrast is used to contrast different meta-path views, while inter-contrast is used to contrast between network schema and meta-path views. We summarize the contributions of this paper as follows:
@@
  • We propose a heterogeneous graph generative learning enhanced contrastive paradigm, leveraging a generative masked autoencoder as a tool for contrastive view augmentation. This represents a meaningful exploration for graph generative adversarial networks.
@@
  • We introduce hierarchical contrastive learning as an enhanced discriminator, combined with novel sampling strategies to align with the generative model. All these modules can better mine self-supervised signals under an end-to-end training scenario.
@@
  • We conduct comprehensive experiments to evaluate GC-HGNN with seventeen baselines methods and eight real datasets. The experimental results show that GC-HGNN is superior to other methods, and is able to generate enhanced views and effectively extract heterogeneous information.
@@ 2.1. Heterogeneous Graph Neural Networks
In recent years, Graph Neural Networks (GNNs) (You et al., 2020; Hu et al., 2020b; Hou et al., 2022; Velickovic et al., 2017; He et al., 2020a) have achieved significant success in fields such as social network analysis and recommender systems. This success relies on their effective modeling of graph structure information and local and global information integration. These graph structures often have different types of nodes and edges. Heterogeneous graph neural networks are designed to better handle heterogeneous graph data.
The currently popular heterogeneous graph neural network models (Wang et al., 2019, 2021; Zhang et al., 2019; Fu et al., 2020; Yu et al., 2022b) can be classified into two categories: meta-path- based methods and types-based methods. The classic meta-path-based models include HAN (Wang et al., 2019), HetGNN (Zhang et al., 2019), and MAGNN (Fu et al., 2020). Specifically, HAN employs a dual-layer attention mechanism at both the node and semantic levels to adaptively capture the relationships between nodes and meta-paths. HetGNN introduces a heterogeneous graph attention propagation mechanism, combining meta-path information and attention mechanisms to learn node representations. MAGNN integrates meta-paths and adaptive mechanisms to learn node representations in heterogeneous graphs by learning meta-path attention weights at both the node and global levels. The type-based models include RGCN (Schlichtkrull et al., 2018), HGT (Hu et al., 2020a), and RHGNN (Yu et al., 2022b). RGCN captures rich semantic relationships by using different weight matrices for different types of edges. HGT achieves efficient representation learning and information propagation on heterogeneous graph data by introducing a multi-head attention mechanism and a heterogeneous graph structure encoder. RHGNN utilizes dedicated graph convolution components and cross-relation message passing modules to learn node representations on heterogeneous graphs. However, the above-mentioned methods only investigate supervised heterogeneous graph models. In order to more deeply explore the potential relationships and patterns in the data, and to effectively utilize unlabeled data, an increasing number of studies are beginning to focus on self-supervised learning methods.
@@ 2.2. Self-supervised Graph Learning
With the rapid development of self-supervised learning in computer vision (He et al., 2020b; Chen et al., 2020) and natural language processing (Devlin et al., 2018; Chowdhery et al., 2023) fields, graph neural networks have drawn inspiration from this progress. Self-supervised graph neural networks often train without labels by generating contrastive views or reconstructing node attributes, thus better leveraging the rich structural information in graph data. Transitioning from general graphs to heterogeneous graphs involves assigning different types or attributes to nodes and edges in the graph, making the graph structure more complex and diverse in representation.
Self-supervised heterogeneous graph neural networks can be categorized into: contrastive and generative. Contrastive learning models (Ren et al., 2019; Wang et al., 2021, 2023) generate enhanced contrastive views and attempt to maximize the mutual information between these views to produce supervised signals. Specifically, HDGI (Ren et al., 2019) utilizes graph convolutional modules and semantic-level attention to learn node representations and maximize local-global mutual information. HeCo (Wang et al., 2021) complements across perspectives through network schema and meta-path views, establishing a sampling contrast method between the two views to achieve self-supervised enhancement. HGCML (Wang et al., 2023) solely utilizes the rich semantic relationships brought by meta-paths for graph enhancement, introducing multi-view contrastive learning. It also proposes sampling strategies for positive samples in both topology and semantics to alleviate sampling bias. Generative models (Zheng et al., 2021; Tian et al., 2023) commonly adopt graph autoencoder and attempt to reconstruct features to obtain supervision. For example, MvDGAE (Zheng et al., 2021) utilizes random edge dropout to enhance views and utilizes autoencoders to enforce the recovery of views for more robust representations. Meanwhile, HGMAE (Tian et al., 2023) further introduces dynamic masking mechanism and position-aware encoding to enhance the model’s performance and utilizes attribute and view dual-recovery to improve denoising capability.
@@ 3. PRELIMINARIES
Meta-path: In an $\operatorname{HIN}G=(V,E,I,J,\phi,\varphi)$, a meta-path $\rho$ is represented as $I_{1}\stackrel{{\scriptstyle J_{1}}}{{\longrightarrow}}I_{2}\stackrel{{\scriptstyle J_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle J_{l}}}{{\longrightarrow}}I_{l+1}$, describing a composite connection between $v_{1}$ and $v_{l}$. Taking Figure 2 as an instance, multiple meta-paths, like “Paper-Author-Paper (PAP)” and “Paper-Subject-Paper (PSP)” can link entities within the HIN. Commonly, different meta-paths can reveal the different inter-dependency information of two movies. The path “PAP” indicates that two papers were written by the same author, while “PSP” illustrates that two papers are of the same subject.
Figure 3. The training framework of our proposed GC-HGNN. (a) The model respectively takes the entire heterogeneous graph and the views based on meta-paths, dividing them into inputs for the network schema View and the meta-path view; (b) the network schema view employs an attention mechanism to focus on one-hop neighbors of each node; (c) the meta-path view utilizes a masked autoencoder to generate views for intra-contrast; (d) the hierarchical contrast uses a proposed dynamic sampling mechanism to generate more hard negative samples to enhance the discriminator.
@@ 4.1. Network Schema View
This section typically takes the entire heterogeneous information network as input, as shown in Figure 3 (a). First, we consider each node in the network as an anchor, forming a subnetwork that includes all the one-hop neighbor nodes. For example, in Figure 3 (b), the subgraph indicated by the circles represents the subnetwork composed of the target node. Next, we put such subnetworks into a heterogeneous neighbor aggregation network, utilizing attention mechanisms to obtain node embeddings. Finally, we select the all embeddings of the desired target node for downstream tasks. Specific implementation process: We project all nodes into the same node space ($\boldsymbol{P}_{r}$ is projection matrix) because the subnetworks contain different node types. We define $e$ as the current node computing the attention score and $\boldsymbol{h}_{e}^{0}$ as its initial embedding. Next, we learn specific aggregation functions for different types of edges to compute attention scores:

### v2

# Generative-Enhanced Heterogeneous Graph Contrastive Learning

@@ Abstract.
Heterogeneous Graphs (HGs) can effectively model complex relationships in the real world by multi-type nodes and edges. In recent years, inspired by self-supervised learning, contrastive Heterogeneous Graphs Neural Networks (HGNNs) have shown great potential by utilizing data augmentation and contrastive discriminators for downstream tasks. However, data augmentation is still limited due to the graph data’s integrity. Furthermore, the contrastive discriminators remain sampling bias and lack local heterogeneous information. To tackle the above limitations, we propose a novel Generative-Enhanced Heterogeneous Graph Contrastive Learning (GHGCL). Specifically, we first propose a heterogeneous graph generative learning enhanced contrastive paradigm. This paradigm includes: 1) A contrastive view augmentation strategy by using a masked autoencoder. 2) Position-aware and semantics-aware positive sample sampling strategy for generating hard negative samples. 3) A hierarchical contrastive learning strategy for capturing local and global information. Furthermore, the hierarchical contrastive learning and sampling strategies aim to constitute an enhanced contrastive discriminator under the generative-contrastive perspective. Finally, we compare our model with seventeen baselines on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on node classification and link prediction tasks.
@@ 1. INTRODUCTION
Figure 1. (a)Heterogeneous graph self-supervised learning with contrastive paradigms; (b)Heterogeneous graph self-supervised learning with generative-contrastive paradigms.
@@
Self-supervised learning (SSL) has been widely applied in the fields of computer vision (CV) (He et al., 2020b; Chen et al., 2020; Grill et al., 2020) and natural language processing (NLP) (Chowdhery et al., 2023; Kenton and Toutanova, 2019), leveraging unlabeled data to learn informative representations. Inspired by this trend, researchers explore the combination of HGNNs and self-supervised learning. As a major paradigm, contrastive HGNNs (Wang et al., 2021; Zhu et al., 2022; Wang et al., 2023; Chen et al., 2023) based SSL have shown great potential because they are more suitable for downstream tasks by setting contrastive view and discriminator. Data augmentation (Wu et al., 2021; Yang et al., 2023) is typically used as a tool to construct contrastive views, as it maximizes mutual information by aligning two views. However, due to the need to ensure graph data’s integrity and unique features, it is difficult to provide reasonable data augmentation to leverage the advantages of contrastive learning (He et al., 2020b; Liu et al., 2021).
@@
Generative learning (Chowdhery et al., 2023; Tian et al., 2023; Hou et al., 2022; Kenton and Toutanova, 2019; Liu et al., 2021) can effectively solve the limitation because it can reconstruct the data distribution without considering downstream tasks and without data augmentation, allowing each node embedding to be enhanced. This idea also provides a foundation for us to propose a generative learning enhanced contrastive paradigm for heterogeneous graphs. As shown in Figure 1, first, HGNNs utilize a meta-path to obtain a graph containing nodes of the same type. Contrastive learning (Wang et al., 2023; Zhu et al., 2022) generates supervised signals by aligning two similar views constructed through random edge dropping; while generative-contrastive learning (Goodfellow et al., 2020; Yang et al., 2023; Wu et al., 2023) reconstructs the intrinsic structures and features of data by training an autoencoder and uses a contrastive discriminator to distinguish them in the embedding space. Despite the obvious advantages of the generative-contrastive paradigm, there are still some challenges for heterogeneous graphs:
@@
CH1: Unbalanced data augmentation. The current augmentation strategies (Wang et al., 2023; Zhu et al., 2022; Lin et al., 2022; Wu et al., 2021), such as structural and feature augmentation, still have limitations. Structural augmentation (Wang et al., 2023; Zhu et al., 2022; Wu et al., 2021), which involves randomly dropout nodes and edges, could disrupt the graph integrity since nodes and edges are the fundamental properties of a graph. On the other hand, feature augmentation (Wang et al., 2021; Lin et al., 2022), which perturbs node attributes, ignores the unique features of each node’s input. In such methods, some nodes have more edges and embeddings than others, resulting in unbalanced supervised signals. Exploring how to better utilize meta-paths to design enhanced contrastive views remains a challenge that needs to be addressed.
@@
CH2: Negative sampling bias. In heterogeneous graphs, contrastive learning (Wang et al., 2023, 2021; Zhu et al., 2022) heavy relies on high-quality negative samples. Moreover, the contrastive discriminator has stricter requirements for sampling under the generative-contrastive paradigm (Wu et al., 2023). Existing sampling strategies (Wang et al., 2023, 2021; Zhu et al., 2022) are usually based on randomness or meta-paths, which often increase false negative samples. Some recent works (Kalantidis et al., 2020; Liu et al., 2023) aim to select or synthesize hard negatives to mitigate the impact of false negatives. However, they are not suitable for heterogeneous graphs and offer limited benefits. Therefore, designing an enhanced positive sampling strategy for heterogeneous graphs is urgently needed.
@@
CH3: Lack local heterogeneous information. Existing methods (Wang et al., 2023; Zhu et al., 2022; Tian et al., 2023) rely solely on meta-paths to capture rich semantic information in heterogeneous graphs and often overlook the influence of one-hop neighbors on target nodes. Network schema leverages direct connections between nodes to capture the local structure of target nodes. In contrast, meta-paths extract higher-order structures. Designing an effective self-supervised mechanism that considers both local and global information can better help the model work.
@@
To tackle the three challenges mentioned above, we propose a heterogeneous graph self-supervised model called Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). For CH1, GC-HGNN employs a generative masked autoencoder (MAE) to enhance contrastive views. The MAE can reconstruct node embeddings without altering the original graph structure and features. The masking strategy averages the edges of each node to ensure that each node receives the same supervised signals during the reconstruction process. For CH2, the model proposes a sampling strategy that combines location-awareness and semantic-awareness. This approach adaptively samples nodes based on the rich semantics unique to heterogeneous graphs and the global positions of nodes in the heterogeneous graph. Finally, for CH3, GC-HGNN utilizes hierarchical contrastive (HCL) learning to capture one-hop and higher-order neighbors. Intra-contrast is used to contrast different meta-path views, while inter-contrast is used to contrast between network schema and meta-path views. We summarize the contributions of this paper as follows:
@@
  • We propose a heterogeneous graph generative learning enhanced contrastive paradigm, leveraging a generative masked autoencoder as a tool for contrastive view augmentation. It is a meaningful exploration for graph generative-contrastive networks.
@@
  • We introduce hierarchical contrastive learning as an enhanced contrastive discriminator, combined with novel sampling strategies to align with the generative model. These modules can better mine self-supervised signals under an end-to-end training scenario.
@@
  • We conduct comprehensive experiments to evaluate GC-HGNN with seventeen baseline methods and eight real datasets. The experimental results show that GC-HGNN is superior to other methods and effectively extracts heterogeneous information.
@@ 2.1. Heterogeneous Graph Neural Networks
In recent years, Graph Neural Networks (GNNs) (You et al., 2020; Hu et al., 2020b; Hou et al., 2022; Veličković et al., 2018; He et al., 2020a) have achieved significant success in fields such as social network analysis and recommender systems. This success relies on their effective modeling of graph structure information and local and global information integration. These graph structures often have different types of nodes and edges. Heterogeneous graph neural networks are designed to better handle heterogeneous graph data.
The currently popular heterogeneous graph neural network models (Wang et al., 2019, 2021; Zhang et al., 2019; Fu et al., 2020; Yu et al., 2022b) can be classified into meta-path-based methods and types-based methods. The classic meta-path-based models include HAN (Wang et al., 2019), HetGNN (Zhang et al., 2019), and MAGNN (Fu et al., 2020). Specifically, HAN employs a dual-layer attention mechanism at both the node and semantic levels to adaptively capture the relationships between nodes and meta-paths. HetGNN introduces a heterogeneous graph attention propagation mechanism, combining meta-path information and attention mechanisms to learn node representations. MAGNN integrates meta-paths and adaptive mechanisms to learn node representations in heterogeneous graphs by learning meta-path attention weights at both the node and global levels. The type-based models include RGCN (Schlichtkrull et al., 2018), HGT (Hu et al., 2020a), and RHGNN (Yu et al., 2022b). RGCN captures rich semantic relationships by using different weight matrices for different types of edges. HGT achieves efficient representation learning and information propagation on heterogeneous graph data by introducing a multi-head attention mechanism and a heterogeneous graph structure encoder. RHGNN utilizes dedicated graph convolution components and cross-relation passing modules to learn node representations on heterogeneous graphs. However, the methods mentioned above only investigate supervised heterogeneous graph models. In order to deeply explore the potential relationships and patterns in the data and to effectively utilize unlabeled data, an increasing number of studies are beginning to focus on self-supervised learning methods.
@@ 2.2. Self-supervised Graph Learning
With the rapid development of self-supervised learning in computer vision (He et al., 2020b; Chen et al., 2020) and natural language processing (Kenton and Toutanova, 2019; Chowdhery et al., 2023) fields, graph neural networks have drawn inspiration from this progress. Self-supervised graph neural networks often train without labels by generating contrastive views or reconstructing node attributes, thus better leveraging the rich structural information in graph data. Transitioning from general graphs to heterogeneous graphs involves assigning different types or attributes to nodes and edges in the graph, making the graph structure more complex and diverse in representation.
Self-supervised heterogeneous graph neural networks can be categorized into contrastive and generative. Contrastive learning models (Ren and Liu, 2020; Wang et al., 2021, 2023) generate enhanced contrastive views and attempt to maximize the mutual information between these views to produce supervised signals. Specifically, HDGI (Ren and Liu, 2020) utilizes graph convolutional modules and semantic-level attention to learn node representations and maximize local-global mutual information. HeCo (Wang et al., 2021) complements across perspectives through network schema and meta-path views, establishing a sampling contrast method between the two views to achieve self-supervised enhancement. HGCML (Wang et al., 2023) solely utilizes the rich semantic relationships brought by meta-paths for graph enhancement, introducing multi-view contrastive learning. It also proposes sampling strategies for positive samples in both topology and semantics to alleviate sampling bias. Generative models (Zheng et al., 2021; Tian et al., 2023) commonly adopt graph autoencoder and attempt to reconstruct features to obtain supervision. For example, MvDGAE (Zheng et al., 2021) utilizes random edge dropout to enhance views and autoencoders to enforce the recovery of views for more robust representations. Meanwhile, HGMAE (Tian et al., 2023) further introduces a dynamic masking mechanism and position-aware encoding to enhance the model’s performance and utilizes attribute and view dual-recovery to improve denoising capability.
@@ 3. PRELIMINARIES
Meta-path: In an $\operatorname{HIN}G=(V,E,I,J,\phi,\varphi)$, a meta-path $\rho$ is represented as $I_{1}\stackrel{{\scriptstyle J_{1}}}{{\longrightarrow}}I_{2}\stackrel{{\scriptstyle J_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle J_{l}}}{{\longrightarrow}}I_{l+1}$, describing a composite connection between $I_{1}$ and $I_{l+1}$. Taking Figure 2 as an instance, multiple meta-paths, like “Paper-Author-Paper (PAP)” and “Paper-Subject-Paper (PSP)” can link entities within the HIN. Commonly, different meta-paths can reveal the different inter-dependency information of two papers. The path “PAP” indicates that two papers were written by the same author, while “PSP” illustrates that two papers are of the same subject.
Figure 3. The training framework of our proposed GC-HGNN. (a) The model respectively takes the entire heterogeneous graph and the views based on meta-paths, dividing them into inputs for the network schema View and the meta-path view; (b) the network schema view employs an attention mechanism to focus on one-hop neighbors of each node; (c) the meta-path view utilizes a masked autoencoder to generate views for intra-contrast; (d) the hierarchical contrast uses a proposed dynamic sampling mechanism to generate more hard negative samples to enhance the contrastive discriminator.
@@ 4.1. Network Schema View
This section typically takes the entire heterogeneous information network as input, as shown in Figure 3 (a). First, we consider each node in the network as an anchor, forming a subnetwork that includes all the one-hop neighbor nodes. For example, in Figure 3 (b), the subgraph indicated by the circles represents the subnetwork composed of the target node. Next, we put such subnetworks into a heterogeneous neighbor aggregation network, utilizing attention mechanisms to obtain node embeddings. Finally, we select all the embeddings of the desired target node for downstream tasks. Specific implementation process: We project all nodes into the same node space ($\boldsymbol{P}_{r}$ is projection matrix) because the subnetworks contain different node types. We define $e$ as the current node computing the attention score and $\boldsymbol{h}_{e}^{0}$ as its initial embedding. Next, we learn specific aggregation functions for different types of edges to compute attention scores: